OpenStudy (anonymous):

There are two consecutive odd numbers such that twice the smaller number plus one is ten more than the greater number. What is the smaller number? 7 9 11 13

OpenStudy (anonymous):

check

OpenStudy (anonymous):

\[2*7=14\] no, since 14 is not ten more than 9

OpenStudy (anonymous):

\[2\times 9=18\] nope

OpenStudy (anonymous):

\[2\times 11=22\]

OpenStudy (anonymous):

a,b is the number for odd number is 2n-1 a is 2n-1 and b is 2(n+1)-1 2a+1=10+b so n=6 hence a=11 b=13

OpenStudy (anonymous):

add 1 get 23. so that is the one!

OpenStudy (anonymous):

so 11